OpenStudy (anonymous):

find the coordinates of p(1,2) with reference to the new axes, when the axes are rotated through an angle of 30 degrees

ganeshie8 (ganeshie8):

use the formula ?

ganeshie8 (ganeshie8):

new coordinates after rotating axes : \(x' = x . cos\theta + y . sin\theta \) \(y' = -x . sin\theta + y . cos\theta \)

OpenStudy (anonymous):

i dint get u